The amino acid methionine, which is essential in human diets, has the molecular formula C5H11NO2S. Determine (a) its molecular mass; (b) the number of moles of H atoms per mole of methionine; (c) number of grams of C per mole of methionine; (d) the number of C atoms in 9.07 mol methionine


1
Expert's answer
2022-03-17T14:58:17-0400

Solution:

(a):

Calculate the molecular mass of methionine (C5H11NO2S):

Mr(C5H11NO2S) = 5×Ar(C) + 11×Ar(H) + Ar(N) + 2×Ar(O) + Ar(S)

Mr(C5H11NO2S) = 5×12.0107 + 11×1.00784 + 14.0067 + 2×15.999 + 32.065 = 149.209 = 149.21

The molecular mass of methionine is 149.21 g/mol


(b):

Calculate the number of moles of H atoms per mole of methionine:

C5H11NO2S → 11H

Thus, 1 mol of methionine contains 11 mol of H atoms

11 moles of H atoms per mole of methionine


(c):

Calculate the number of grams of C per mole of methionine:

C5H11NO2S → 5C

Thus, 1 mol of methionine contains 5 mol of C atoms

The molar mass of carbon (C) is 12.0107 g/mol

Therefore,

Mass of C = (5 mol C) × (12.0107 g C / 1 mol C) = 60.0535 g C = 60.05 g C

60.05 grams of C per mole of methionine


(d):

Calculate the number of C atoms in 9.07 mol methionine

1 mol of methionine contains 5 mol of C atoms

Thus, 9.07 mol of methionine contains 5×9.07 = 45.35 mol of C atoms

One mole of any substance contains 6.022×1023 particles

Therefore,

Number of C atoms = (45.35 mol) × (6.022×1023 atoms / 1 mol) = 2.73×1025 atoms

2.73×1025 carbon atoms in 9.07 mol methionine